A car is driving at 25 m/s and sees a red light. The car comes to stop over a distance of 100 m. What is their acceleration? Time?


1
Expert's answer
2021-02-06T10:06:46-0500

"v=\sqrt{2as}, ~a=\frac{v^2}{2s}=3.125~\frac{m}{s^2},"

"t=\frac va=8~s."